# Precise calculation of the EFT likelihood with primordial non-Gaussianities

### v1

@@ Abstract
We perform a precise calculation of the EFT conditional likelihood for large-scale structure (LSS) via the saddle-point expansion method in the presence of primordial non-Gaussianities (PNG). The precision is manifested at two levels: one corresponding to the consideration of higher-order noise terms, and the other to the inclusion of contributions around the saddle points. In computing the latter, we encounter the same issue of the negative modes as in the context of false vacuum decay, which necessitates deforming the original integration contour into the combination of the steepest descent contours to ensure a convergent and real result. We demonstrate through detailed calculations that, upon the incorporation of leading-order PNG, both types of extensions introduce irreducible field-dependent contributions to the conditional likelihood. This insight motivates the systematic inclusion of additional effective terms within the forward modeling framework. Our work facilitates Bayesian forward modeling under non-Gaussian initial conditions, thereby enabling more stringent constraints on the parameters describing PNG.
@@ 1 Introduction
Primordial non-Gaussianities (PNG) are a pivotal and enduring area of interest in contemporary cosmology. On the one hand, they constitute a fundamental prediction of inflationary theory [1, 2, 3, 4, 5, 6], potentially providing a crucial observational window into the relics of the early universe (e.g. physics beyond the standard model [7, 8]). On the other hand, non-Gaussian initial conditions can also impact the evolution of the dark matter and galaxy bias [9, 10, 11, 12, 13], consequently leaving characteristic imprints on the large-scale structure (LSS) of the universe. Recent studies have been endeavoring to capture traces of PNG from both theoretical [14, 15, 16] and experimental [17, 18, 19] perspectives.
@@
This motivates the continuous development of more advanced methodologies to constrain the parameters describing PNG, among which PNG Bayesian forward modeling [20, 21] emerges as a particularly prominent approach, owing to its capacity to optimally exploit the available information. In the spirit of this approach, it is essential to account for various effects within the context of specified initial conditions, directly predict the present-day galaxy distribution, and subsequently compare the results with observational data. Within this framework, a key quantity requiring a semi-analytical treatment is the conditional likelihood $\mathcal{P}[\delta_{g}|\delta]$, which is the functional probability of observing a galaxy field $\delta_{g}$ given a realization of the matter field $\delta$. Currently, likelihood computation methods can be broadly classified into two main categories: the traditional EFT method [22] and the functional method [23, 24].
To provide more reliable predictions, we generally seek to incorporate as many cosmological effects as possible in the forward modeling. Among these, the non-ideal fluid effects induced by small-scale fluctuations on the large-scale universe, predicted by the Effective Field Theory of LSS [^1] The extension of this theory to non-Gaussian initial conditions can be found in [25, 26] (EFTofLSS) [27, 28, 29, 30, 31, 32, 33], are particularly noteworthy. We will hereafter refer to the likelihood incorporating this effect as the EFT likelihood, abbreviated simply as likelihood when context permits. Therefore, we are more inclined toward the functional approach due to its enhanced flexibility and universality. Specifically, this approach enables the direct derivation of the likelihood expression by performing a functional Fourier transform of the LSS partition function, in which different cosmological effects are systematically encoded as corresponding vertices within the formalism. Consequently, during the simulation, we are able to directly employ the conditional likelihood after effectively encapsulating the contributions of these equivalent terms, thereby accounting for the associated physical effects. The functional approach for calculating the EFT likelihood was first proposed in [23], where the authors approximated the functional integral by incorporating contributions from the saddle points and identified terms that were neglected by the traditional approach. In our previous work [24], we developed the functional approach and demonstrated that this functional integral can be reformulated as a path integral in three-dimensional Euclidean space. This reformulation allows for the application of the saddle-point expansion method, which was originally applied in quantum mechanics and quantum field theory for computing the tunneling rates [34, 35].
@@
The present work is devoted to further developing this method by performing a precise calculation of the conditional likelihood in the presence of PNG. The standard approach to incorporating PNG [36, 26] introduces an effective vertex cubic in the initial matter field within the partition function [37]. In [23], the authors investigated the case in which only the leading-order noise contributions are included, and argued that under this approximation, the corrections from PNG to both the matter likelihood and the joint likelihood are identical, thereby yielding a conditional likelihood that is consistent with that obtained under Gaussian initial conditions. Nevertheless, this work establishes that once more precise calculations are taken into account, nontrivial corrections to the conditional likelihood inevitably emerge.
@@ 2 Review of the saddle-point expansion method
For the sake of clarity and familiarity, in this section we present a brief review of the saddle-point expansion method for computing the EFT likelihood, illustrated via a concrete example. This method was initially proposed in quantum mechanics and quantum field theory for the calculation of decay rates, leading to reliable results that are consistent with the quantum mechanical estimates of tunneling rates [34, 35]. More recently, this method has been applied to the path-integral approach to the LSS and yields more precise calculation of the conditional likelihood [24]. Specifically, we will focus on the origin of the negative (complex) modes and demonstrate how the original, potentially ill-defined integration contour can be systematically deformed into a sum of the steepest descent contours to restore convergence and physical consistency.
Suppose now we aim to calculate the joint likelihood of the following form (identical to that considered in [24, 23]):
@@
In this example, there is only one saddle-point solution. The contribution from the saddle point will dominate the path integral, as it corresponds to the region where the action exhibits minimal oscillatory behavior. Thus we can decompose the field around the saddle point as $\boldsymbol{\phi}_{g}=\bar{\boldsymbol{\phi}}_{g}+\boldsymbol{\varphi}_{g}$, where $\boldsymbol{\varphi}_{g}$ is the higher-order fluctuations  [^2] In the context of quantum field theory, the field expansion should be $\phi=\bar{\phi}+\hbar^{1/2}\varphi$ (e.g. see [42]), so the expansion term can be systematically treated as a perturbative variable. Moreover, it has been illustrated in [24] that an analogous result holds at large scales of the universe, thereby justifying the applicability of this decomposition in our analysis. (also referred to as the quantum fluctuation in quantum theory). Substituting this decomposition into eq. (2.1), we can obtain
@@
It can be seen that the saddle-point contribution $S_{g}[\bar{\boldsymbol{\phi}}_{g}]$ is independent of the remaining integration and may therefore be extracted as an overall prefactor. As for the fluctuations around the saddle point, one must compute the eigenvalues of the second derivative of the action at the saddle point, then apply the Gaussian integral formula to evaluate their contributions. However, a complication arises when negative (complex) eigenvalues of $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$ appear in the calculation, rendering the entire integral ill-defined, i.e. leading to unphysical imaginary results [35]. For instance, from the expression for $\mathcal{M}_{g}$ in eq. (2.3), we can immediately observe that it is a general three-dimensional matrix. Consequently, its eigenvalues may involve negative or imaginary parts. We can examine this through direct calculations. Incorporating these eigenvalues into our framework would yield a complex likelihood, which is incompatible with its physical interpretation as a probability.
In fact, the presence of the negative modes is not an isolated case, but rather a longstanding issue that has persisted in the computation of tunneling rates. In the context of false vacuum decay, the predecessors ultimately identified the steepest descent method within the framework of Picard-Lefschetz theory as a solution to resolve this problem, as illustrated in [35, 39, 40]. Also, this approach has more recently been demonstrated to be directly transplanted to the computation of the likelihood, yielding robust and reliable results [24]. Following the insights developed in the earlier studies, we summarize the overall strategy in the following steps:
@@
This implies that we extend the field $\boldsymbol{\phi}_{g}$ to the complex plane and perform the path integral on a middle-dimensional contour [39]. In our example, it can be seen in eq. (2.4) that this complexification does not introduce any new saddle-point solutions, so we can retain the previous solution.
@@
To achieve this, we parameterize the family of contours passing through the saddle point $\bar{\boldsymbol{\phi}}_{g}$ via $u$, and impose the boundary condition $\boldsymbol{\phi}_{g}(\boldsymbol{k},u\rightarrow-\infty)=\bar{\boldsymbol{\phi}}_{g}$, which implies in the limit of $u\rightarrow-\infty$, we recover the saddle-point solution. To ensure the steepest descent of the real part of $S_{g}$, we choose the holomorphic function as $\mathcal{I}[\boldsymbol{\phi}_{g}]=-S_{g}[\boldsymbol{\phi}_{g}]$, and then define the Morse function $f[\boldsymbol{\phi}_{g}]={\rm Re}(\mathcal{I}[\boldsymbol{\phi}_{g}])$. Thus, the gradient flow equation for $\bar{\boldsymbol{\phi}}_{g}$ is [41]
$\frac{\partial\boldsymbol{\phi}_{g}(\boldsymbol{k},u)}{\partial u}=-\overline{\left(\frac{\delta\mathcal{I}[\boldsymbol{\phi}_{g}(\boldsymbol{k},u)]}{\delta\boldsymbol{\phi}_{g}(\boldsymbol{k},u)}\right)},\ \ \ \ {\frac{\partial\overline{\boldsymbol{\phi}_{g}(\boldsymbol{k},u)}}{\partial u}}=-\left(\frac{\delta\mathcal{I}[\boldsymbol{\phi}_{g}(\boldsymbol{k},u)]}{\delta\boldsymbol{\phi}_{g}(\boldsymbol{k},u)}\right),$ (2.6)
note that the overline here does not denote a saddle point, but rather the complex conjecture (following the convention adopted in [39]). By examining the non-positivity of $\partial f/\partial u$, we can confirm that the condition imposed by the equation indeed selects the desired steepest descent contour. After substituting the expression for the action eq. (2.2) and employing the classical equation of motion eq. (2.4), we arrive at
@@
The kernel of the steepest descent method lies in approximating the original integration contour by the combination of the steepest descent contours that is homologous to it  [^3] This procedure is generally feasible, as all convergent integration contours within this manifold — and their equivalence relations — form a homology group, see [24] for details., thereby ensuring a convergent and real result. In this way, the calculation of the likelihood is reformulated as evaluating the path integral over all contours in the complex plane constrained by the given equation eq. (2.7).
iii) Following the above spirit, each Lefshcetz thimble terminates at the convergent regions of the integral or another saddle point. Therefore, each thimble provides a convergent path integral along the complex integration contour in field space.
This conclusion is guaranteed by the Picard-Lefschetz theory [40, 41], and can be explicitly verified within our example. A separation of variables approach can be applied to eq. (2.7) upon adopting the following ansatz: $\boldsymbol{\phi}_{g}(\boldsymbol{k},u)=\sum_{n}g_{n}(u)\chi_{n}(\boldsymbol{k})$, where $g_{n}(u)\in\mathbb{R}$ and the subscript “$n$” distinguishes different directions. After implementing this ansatz, we obtain an eigenequation involving $g_{n}(u)$ and $\chi_{n}(\boldsymbol{k})$
@@
here, we have introduced the real eigenvalues $m_{n}$, since the right-hand side (r.h.s) of this equation is manifestly real. Our task now reduces to solving the corresponding eigenequation
@@
From the explicit form of $\mathcal{M}_{g}$ in eq. (2.3), it is evident that $\mathcal{M}_{g}^{*}$ is a general three-dimensional matrix, which renders the eigenvalue problem analytically challenging. Moreover, since $\mathcal{M}_{g}^{*}$ is neither Hermitian nor anti-Hermitian, we cannot generally expect its eigenstates to be pure real or imaginary. To address this issue, we consider this eigenequation eq. (2.9) together with its complex conjecture $\mathcal{M}_{g}\,\chi_{n}(\boldsymbol{k})=m_{n}\,\chi_{n}^{*}(\boldsymbol{k})$, and construct from them a Hermitian eigenequation
@@
This equation is now amenable to solution, with all eigenvalues $m_{n}$ being real by construction. On the other hand, from the eigenequation on the r.h.s of eq. (2.8), it is evident that the solution for $g_{n}(u)$ takes the form $g_{n}(u)\sim a_{n}\,\exp\,(m_{n}u)$, where $a_{n}\in\mathbb{R}$. Under the boundary condition we have imposed, $g_{n}(u\rightarrow-\infty)=0$, it can be shown that all eigenvalues $m_{n}$ are positive and real, which ensures both the convergence and reality of the integral along the Lefschetz thimbles.
iv) With these thimbles in field space, we can approximate the integration contour as the combination of different thimbles fulfilling the conditions to obtain the real result.
@@
Here, we adopt a more rigorous formulation  [^4] In fact, eq. (2.10) admits six eigenvalues, but only three are of interest in our context. Due to the algebraic approximation took in the calculation, the result of $\sum_{n}m_{n}$ could be either $P_{\epsilon_{g}}-P_{\rm in}^{-1}$ or $-P_{\epsilon_{g}}+P_{\rm in}^{-1}$ (we need to select the non-negative solution). Consequently, we express the result in terms of its absolute value which is non-negative.. Plugging this into eq. (2.5) allows us to compute the likelihood
@@
Let us provide a detailed account of the steps involved in the calculation: in the first step, we approximate the integration contour as the combination of different Lefschetz thimbles $C$, in accordance with the implementation of the steepest descent method. Next, we substitute the expression for $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$. In the second line, we first employ the gradient flow equation to constrain the integration contour to be composed of all steepest descent contours in the complex plane. Finally, we make use of eq. (2.11).
In summary, the saddle-point expansion method approximates the full path integral by incorporating contributions from the saddle points and their neighborhoods, which may, in certain cases, yields complex results. However, when employing the steepest descent method, although the integrals along individual Lefschetz thimbles may lead to imaginary contributions, these imaginary parts are exactly canceled by those from other thimbles, resulting in a manifestly real total expression. The cancellation of these imaginary components and the guarantee of the real result constitute the kernel spirit of the application of the Picard-Lefschetz theory. This method has demonstrated remarkable efficacy in calculating the non-perturbative effect of quantum tunneling, which facilitates the estimation of the temporal occurrence of a first-order phase transition in the early universe (see for example the details in [43, 44] and application in [45]). We also anticipate that this method could prove equally effective in the calculation of other Euclidean-space path integrals, thereby aiding in more precise calculations of the corresponding physical quantities. In fact, the use of the term “precise” in the title of this paper is a deliberate inheritance from the terminology employed in [40] to characterize the computation of the tunneling rates. Moreover, the existence of the negative modes does not indicate a flaw in the theory, but rather a consequence of incorrectly applying the steepest descent method, as illustrated in [39].

### v3

# Precision calculation of the EFT likelihood with primordial non-Gaussianities

@@ Abstract
We perform a precision calculation of the effective field theory (EFT) conditional likelihood for large-scale structure (LSS) using the saddle-point expansion method in the presence of primordial non-Gaussianities (PNG). The precision is manifested at two levels: one corresponding to the consideration of higher-order noise terms, and the other to the inclusion of contributions from the saddle-point fluctuations. When computing the latter, we encounter the same issue of the negative modes as in the context of false vacuum decay, which necessitates deforming the original integration contour into a combination of the steepest descent contours to ensure a convergent and real result. We demonstrate through detailed calculations that, upon incorporating leading-order PNG, both types of extensions introduce irreducible field-dependent contributions to the conditional likelihood. This insight motivates the systematic inclusion of additional effective terms within the forward modeling framework. Our work facilitates Bayesian forward modeling under non-Gaussian initial conditions, thereby enabling more stringent constraints on the parameters describing PNG.
@@ 1 Introduction
Primordial non-Gaussianities (PNG) are a pivotal and enduring area of interest in contemporary cosmology. On the one hand, they constitute a fundamental prediction of inflationary theory [1, 2, 3, 4, 5, 6], potentially providing a crucial observational window into the relics of the early universe (e.g. relics from physics beyond the standard model [7, 8]). On the other hand, non-Gaussian initial conditions can also impact the evolution of the dark matter and galaxy bias [9, 10, 11, 12, 13], consequently leaving characteristic imprints on the large-scale structure (LSS) of the universe. Recent studies have been endeavoring to capture traces of PNG at large scales from both theoretical [14, 15, 16] and experimental [17, 18, 19] perspectives.
@@
This motivates the continuous development of more advanced methodologies to constrain the parameters describing PNG, among which PNG Bayesian forward modeling [20, 21] is particularly prominent, as it optimally exploits all available information without the need of summary statistics. In the spirit of this approach, one should first account for various cosmological effects under specified initial conditions, then directly predict the present-day galaxy distribution, and finally compare the results with observational data until convergent. Within this framework, a key quantity requiring a semi-analytical treatment is the conditional likelihood $\mathcal{P}[\delta_{g}|\delta]$, which is the functional probability of observing a galaxy field $\delta_{g}$ given a realization of the matter field $\delta$. In the standard approach, one typically uses the properties of conditional probability to express this as a ratio of two terms: the matter likelihood $\mathcal{P}[\delta]$ and the joint likelihood $\mathcal{P}[\delta,\delta_{g}]$.
To date, considerable effort has been devoted to deriving analytic expressions for the two likelihoods based on perturbation theory [22, 23, 24, 25, 26, 27]. In this framework, the effective field theory of large-scale structure (EFTofLSS) [28, 29, 30, 31, 32, 33, 34] can be incorporated naturally into the likelihood construction. For instance, Schmidt et al. [22] integrate out small-scale degrees of freedom, coarse-grain the underlying dynamics, and thereby obtain an EFT likelihood that governs the statistical properties on large scales. However, when calculations are based on standard perturbation theory (SPT) [35], one will inevitably omit higher-order contributions to the likelihood, as the evolution equation of $n$-point correlation functions involves higher-order information. [^1] This can be found by comparing the results of [22] and [23]. See [24] for a detailed discussion of the underlying reasons. Consequently, we favor the recently developed functional approach [23, 24], which solves this issue. Its origins can be traced to the time-sliced perturbation theory (TSPT) [36, 37, 38], where it was shown that TSPT reproduces results from SPT without requiring higher-order information. For likelihood calculations, this method yields the likelihood directly by performing a functional Fourier transform of the LSS partition function, reducing the problem to a path integral in three-dimensional Euclidean space [24]. In this way, a wide range of cosmological effects can be incorporated phenomenologically by mapping them to distinct terms in the partition function, leading to a likelihood expression that consistently accounts for these effects. [^2] In fact, this advantage of the functional approach extends well beyond likelihood computations. Several studies have employed the “effective likelihoods” as foundational inputs for deriving Feynman rules [38], renormalization group equations [39, 40], and other components of the theoretical framework.
@@
In our previous work [24], we extended this functional approach beyond the saddle-point approximation. By adapting the saddle-point expansion technique, which was developed in quantum field theory to compute tunneling rates [41, 42], we evaluated the additional contributions arising from fluctuations around the saddle points. Building on these results, the present study carries out a precision calculation of the EFT conditional likelihood in the presence of PNG. The standard approach to incorporating PNG [43, 44] introduces an effective vertex cubic in the initial matter field within the partition functions [40]. In [23], the authors investigated the case in which only the leading-order noise contributions are included, and argued that under this approximation, the corrections from PNG to both the matter likelihood and the joint likelihood are identical, thereby yielding a conditional likelihood that is consistent with that obtained under Gaussian initial conditions. Nevertheless, this work establishes that once more precision calculations are taken into account, nontrivial corrections to the conditional likelihood inevitably emerge.
@@ 2 Review of the saddle-point expansion method
For the sake of clarity and familiarity, in this section we present a brief review of the saddle-point expansion method for computing the EFT likelihood, illustrated via a concrete example. This method was initially proposed in quantum mechanics and quantum field theory for the calculation of decay rates, and has led to reliable results that are consistent with the quantum mechanical estimates of tunneling rates [41, 42]. More recently, this method has been applied to the path-integral approach to the LSS, yielding more precision calculation of the conditional likelihood [24]. Specifically, we will focus on the origin of the negative (complex) modes and demonstrate how the original, potentially ill-defined integration contour can be systematically deformed into a sum of the steepest descent contours, thereby restoring convergence and physical consistency.
Suppose now we aim to calculate the joint likelihood of the following form, which is identical to that considered in [24, 23]:
@@
In this example, there is only one saddle-point solution. The contribution from the saddle point will dominate the path integral, as it corresponds to the region where the action exhibits minimal oscillatory behavior. Therefore, we can decompose the field around the saddle point as $\boldsymbol{\phi}_{g}=\bar{\boldsymbol{\phi}}_{g}+\boldsymbol{\varphi}_{g}$, where $\boldsymbol{\varphi}_{g}$ represents the higher-order fluctuation terms, [^3] In the context of quantum field theory, the field expansion typically takes the form $\phi=\bar{\phi}+\hbar^{1/2}\varphi$ (e.g. see [49]). This allows the expansion term to be systematically treated as a higher-order variable. Moreover, it has been demonstrated in [24] that an analogous result holds at large scales of the universe, thereby justifying the applicability of this decomposition in our analysis. which are also referred to as quantum fluctuations in quantum field theory. Substituting this decomposition into eq. (2.1), we can obtain
@@
It can be seen that the saddle-point contribution $S_{g}[\bar{\boldsymbol{\phi}}_{g}]$ is independent of the remaining integration and may therefore be extracted as an overall prefactor. For the fluctuations around the saddle point, one must compute the eigenvalues of the second derivative of the action at the saddle point and then apply the Gaussian integral formula to evaluate their contributions. However, a complication arises when negative (complex) eigenvalues of $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$ appear in the calculation, rendering the entire integral ill-defined, i.e. leading to unphysical imaginary results [42]. For instance, from the expression for $\mathcal{M}_{g}$ in eq. (2.3), we can immediately observe that it is a general three-dimensional matrix. Consequently, its eigenvalues may involve negative or imaginary parts (we can examine this through direct calculations). Incorporating these eigenvalues into our framework would result in a complex likelihood, which is incompatible with its physical interpretation as a probability.
In fact, the presence of negative modes is not an isolated case, but rather a longstanding issue that has persisted in the computation of tunneling rates. In the context of false vacuum decay, the predecessors ultimately identified the steepest descent method within the framework of Picard-Lefschetz theory as a solution to this problem, as illustrated in [42, 46, 47]. This approach has more recently been demonstrated to be directly applicable to the computation of the likelihoods, yielding robust and reliable results [24]. Following the insights developed in earlier studies, we summarize the overall strategy in the following steps:
@@
This implies that we extend the field $\boldsymbol{\phi}_{g}$ to the complex plane and perform the path integral on a middle-dimensional contour [46]. In our example, as seen in eq. (2.4), this complexification does not introduce any new saddle-point solutions. Therefore, we can retain the previous solution.
@@
To achieve this, we parameterize the family of contours passing through the saddle point $\bar{\boldsymbol{\phi}}_{g}$ via $u$, and impose the boundary condition $\boldsymbol{\phi}_{g}(\boldsymbol{k},u\rightarrow-\infty)=\bar{\boldsymbol{\phi}}_{g}$. This implies that in the limit of $u\rightarrow$-$\infty$, we recover the saddle-point solution. To ensure the steepest descent of the real part of $S_{g}$, we choose the holomorphic function as $\mathcal{I}[\boldsymbol{\phi}_{g}]=-S_{g}[\boldsymbol{\phi}_{g}]$, and then define the Morse function $f[\boldsymbol{\phi}_{g}]={\rm Re}(\mathcal{I}[\boldsymbol{\phi}_{g}])$. Thus, the gradient flow equation for $\bar{\boldsymbol{\phi}}_{g}$ is [48]
$\frac{\partial\boldsymbol{\phi}_{g}(\boldsymbol{k},u)}{\partial u}=-\overline{\left(\frac{\delta\mathcal{I}[\boldsymbol{\phi}_{g}(\boldsymbol{k},u)]}{\delta\boldsymbol{\phi}_{g}(\boldsymbol{k},u)}\right)},\ \ \ \ {\frac{\partial\overline{\boldsymbol{\phi}_{g}(\boldsymbol{k},u)}}{\partial u}}=-\left(\frac{\delta\mathcal{I}[\boldsymbol{\phi}_{g}(\boldsymbol{k},u)]}{\delta\boldsymbol{\phi}_{g}(\boldsymbol{k},u)}\right).$ (2.6)
Note that the overline here does not denote a saddle point, but rather the complex conjecture (following the convention adopted in [46]). By examining the non-positivity of $\partial f/\partial u$, we can confirm that the condition imposed by the equation indeed selects the desired steepest descent contour. After substituting the expression for the action from eq. (2.2) and employing the classical equation of motion from eq. (2.4), we arrive at
@@
The kernel of the steepest descent method lies in approximating the original integration contour by a combination of steepest descent contours that is homologous to it, [^4] This procedure is generally feasible, as all convergent integration contours within this manifold – and their equivalence relations – form a homology group. See [24] for details. thereby ensuring a convergent and real result. In this way, the calculation of the likelihood is reformulated as evaluating the path integral over all contours in the complex plane that are constrained by the given equation eq. (2.7).
iii) Following the above spirit, each Lefschetz thimble terminates at the convergent regions of the integral or at another saddle point. Therefore, each thimble provides a convergent path integral along the complex integration contour in field space.
This conclusion is guaranteed by the Picard-Lefschetz theory [47, 48] and can be explicitly verified within our example. A separation of variables approach can be applied to eq. (2.7) by adopting the following ansatz: $\boldsymbol{\phi}_{g}(\boldsymbol{k},u)=\sum_{n}g_{n}(u)\chi_{n}(\boldsymbol{k})$, where $g_{n}(u)\in\mathbb{R}$ and the subscript “nn” distinguishes different directions. After implementing this ansatz, we obtain an eigen-equation involving $g_{n}(u)$ and $\chi_{n}(\boldsymbol{k})$
@@
where we have introduced the real eigenvalues $m_{n}$, since the right-hand side (r.h.s) of this equation is manifestly real. Our task now reduces to solving the corresponding eigen-equation
@@
From the explicit form of $\mathcal{M}_{g}$ in eq. (2.3), it is evident that $\mathcal{M}_{g}^{*}$ is a general three-dimensional matrix, which renders the eigenvalue problem analytically challenging. Moreover, since $\mathcal{M}_{g}^{*}$ is neither Hermitian nor anti-Hermitian, we cannot generally expect its eigenstates to be purely real or imaginary. To address this issue, we consider this eigen-equation eq. (2.9) together with its complex conjecture $\mathcal{M}_{g}\,\chi_{n}(\boldsymbol{k})=m_{n}\,\chi_{n}^{*}(\boldsymbol{k})$, and construct from them a Hermitian eigen-equation
@@
This equation is now amenable to solution, with all eigenvalues $m_{n}$ being real by construction. On the other hand, from the eigen-equation on the r.h.s of eq. (2.8), it is evident that the solution for $g_{n}(u)$ takes the form $g_{n}(u)\sim a_{n}\,\exp\,(m_{n}u)$, where $a_{n}\in\mathbb{R}$. Under the boundary condition we have imposed, $g_{n}(u\rightarrow-\infty)=0$, it can be shown that all eigenvalues $m_{n}$ are positive and real. This ensures both the convergence and reality of the integral along the Lefschetz thimbles.
iv) With these thimbles in field space, we can approximate the integration contour as a combination of different Lefschetz thimbles that homology to the original integral contour to obtain a convergent and real result.
@@
Here, we adopt a more rigorous formulation. [^5] In fact, eq. (2.10) admits six eigenvalues, but only three are of interest in our context. Due to the algebraic approximation taken in the calculation, the result of $\sum_{n}m_{n}$ could be either $P_{\epsilon_{g}}-P_{\rm in}^{-1}$ or $P_{\epsilon_{g}}+P_{\rm in}^{-1}$ (we need to select the non-negative solution). Consequently, we express the result in terms of its absolute value which is non-negative. Plugging this into eq. (2.5) allows us to compute the likelihood
@@
Let us provide a detailed account of the steps involved in the calculation. In the first line, we first approximate the integration contour as the combination of different Lefschetz thimbles $C$, in accordance with the implementation of the steepest descent method. Then we substitute the expression for $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$. In the second line, we first employ the gradient flow equation to constrain the integration contour to be composed of all steepest descent contours in the complex plane. Finally, we make use of eq. (2.11).
In summary, the saddle-point expansion method approximates the full path integral by incorporating contributions from the saddle points and their neighborhoods, which may, in certain cases, yields complex results. However, when employing the steepest descent method, although the integrals along individual Lefschetz thimbles may lead to imaginary contributions, these imaginary parts are exactly canceled by those from other thimbles, resulting in a manifestly real total expression. The cancellation of these imaginary components and the guarantee of a real result constitute the core spirit of the application of the Picard-Lefschetz theory. This method has demonstrated remarkable efficacy in calculating the non-perturbative effects of quantum tunneling, facilitating the estimation of the temporal occurrence of a first-order phase transition in the early universe (see, for example, [50, 51] for details and [52] for applications). We also anticipate that this method could prove equally effective in calculating other Euclidean-space path integrals, thereby aiding in more precision calculations of the corresponding physical quantities. In fact, the use of the term “precision” in the title of this paper is a deliberate inheritance from the terminology employed in [47] to characterize the computation of tunneling rates. Moreover, the existence of negative modes does not indicate a flaw in the theory, but rather a consequence of incorrectly applying the steepest descent method, as illustrated in [46].